OpenStudy (anonymous):

Please help I still don't understand the question please help. Solve the equation 3x-7=5x+ 19

OpenStudy (anonymous):

do it and ill correct it :)

OpenStudy (anonymous):

@Sevier Show your work and we're more than happy to help you out :)

OpenStudy (anonymous):

3x-7=-7 5x+19=19

OpenStudy (anonymous):

@Chlorophyll this is my first time making Alegbra

OpenStudy (anonymous):

3x-7=5x+ 19 put all the x's on one side and all the constants (numbers) on the other side

OpenStudy (anonymous):

Yup, pretend x as the number of ice creams, and the constants as the amount of money Therefore, you need to group them together as HawkCrimson's instruction :)

OpenStudy (anonymous):

do that and ill let u know if it is right :)

OpenStudy (anonymous):

32-7=52+19

OpenStudy (anonymous):

is the question 3x-7=5x+ 19 or 3x-7=-7 5x+19=19???

OpenStudy (anonymous):

3x-7=5x+19

OpenStudy (anonymous):

First which one are X terms here? Put them together :/

OpenStudy (anonymous):

ok so now bring all the x's to the right and all the numbers to the left :)

OpenStudy (anonymous):

I still don't understand the question at all. Please help.